You draw from bag 5 times, and there are 4 red and 6 black balls in the bag. what's the probability that 3 red will be drawn with no replacement?
Answer by ewatrrr(24785)

 
Hi
You draw from bag 5 times that has 4 red and 6 black balls in it, with no replacement
P(3 reds) = (4C3*6C2)/10C5 = 4*15/152 = 60/152 = 15/38